OpenStudy (anonymous):

^3√-27x^6 Carrots represent powers.

OpenStudy (hoblos):

-3x^2

OpenStudy (dumbcow):

\[\sqrt[3]{-27x^{6}} = \sqrt[3]{-27}*\sqrt[3]{x^{6}} = -3x^{2}\]

OpenStudy (dumbcow):

in general \[\large \sqrt[n]{x^{m}}= x^{m/n}\]